Shannon’s Brewery in Keller, Texas is considering using a sales team to help it penetrate market areas not served by its current distributors.  However, Shannon Carter, the CEO, wants to know whether it would be cheaper to hire an independent outside sales force rather than hire and train his own sales personnel.  Estimates are that the costs of hiring, training, and maintaining his own salesforce would entail fixed costs of salaries and administrative overhead of $300,000 per year.  These salespeople would also receive a 3.2% commission on all sales.  If going with the independent sales force option, all costs would be variable in that these reps would be paid a commission of 9.9% on all sales.  At what level of sales will the costs of each sales force be equal?  Round your answer to the nearest one dollar.
Answer #1

Answer:

Let the level of sales be ‘$ X’.
At $ X level, the cost of hiring outside sales force and cost of hiring and training his own sales personnel will be SAME
Cost of hiring independent outside sales force = $ X x 9.90% = 0.099X
Cost of hiring and training own sales force = 300000 + ($ X x 3.2%) = 300000 + 0.032X
Equation: 0.0990X = 300000 + 0.032X
0.099X – 0.032X = 300000
0.067X = 300000
X = $ 44,77,611.94
Answer = $ 44,77,612 will be the level of sales at which cost of each sales force will be equal.